OpenStudy (anonymous):

Hi, what is the equation (2) that being referred to repeatedly in Problem Set 3?

OpenStudy (anonymous):

Do you mean Exercises 2a, 2e, & 2h from section 2B of Problem set 3? If so, then I believe the question is referring to the equations listed in the previous problem on the problem set. Thus, the question is asking you to find the inflection points of the following graphs: \[y = x^3 − 3x + 1\]\[y = x/(x + 4)\]\[y = e^{-x^2}\] Does this help?

OpenStudy (anonymous):

This problem set was originally designed for use in a different calculus course that had some supplemental readings. The reading for approximation is available here: http://ocw.mit.edu/courses/mathematics/18-01-single-variable-calculus-fall-2006/readings/a_approximations.pdf Equation (2) appears in that reading (which is kind of interesting for a slightly different take on approximation than presented in this course).

OpenStudy (anonymous):

Thanks Splash_Dance - it was actually 2A-1 and 3. creeksider, thanks for pointing it out.